In a segment of DNA 3.2 kilobases are present. If DNA segment has 820 adenine molecules, then what will be the bomber of cytosine?

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Chargaff's equivalency rule
(A+G)/(T+C) = 1
A+G = T+C
A = T
G = C
If A = 820 then T = 820
A + T = 1640
G + C = 3200 -1640 = 1560
Hence 
G = C = 780